# Do AI Models Perform Human-like Abstract Reasoning Across Modalities?

### v3

@@ Abstract
OpenAI’s o3-preview reasoning model exceeded human accuracy on the ARC-AGI benchmark, but does that mean state-of-the-art models recognize and reason with the abstractions that the task creators intended? In this work, we investigate the abstraction abilities of AI models using the ConceptARC benchmark. We evaluate models under settings that vary the input modality (textual vs. visual), whether the model is permitted to use external Python tools, and, for reasoning models, the amount of reasoning effort. In addition to measuring output accuracy, we perform fine-grained evaluation of the natural-language rules that models generate to explain their solutions. This dual evaluation allows us to assess whether models solve tasks using the abstractions that ConceptARC was designed to elicit, rather than relying on surface-level patterns. Our results show that, while some models using text-based task representations match human output accuracy, the best models’ rules are frequently based on surface-level “shortcuts”, and capture intended abstractions substantially less often than do humans. Thus their capabilities for general abstract reasoning may be overestimated by evaluations based on accuracy alone. In the visual modality, AI models’ output accuracy drops sharply, yet our rule-level analysis reveals that models might be underestimated, as they still exhibit a substantial share of rules that capture intended abstractions, but are often unable to correctly apply these rules. In short, our results show that models still lag humans in abstract reasoning, and that using accuracy alone to evaluate abstract reasoning on ARC-like tasks may overestimate abstract-reasoning capabilities in textual modalities and underestimate it in visual modalities. We believe that our evaluation framework offers a more faithful picture of multimodal models’ abstract reasoning abilities and a more principled way to track progress toward human-like, abstraction-centered intelligence.
@@ 1 Introduction
The ability to quickly form abstractions and reason with them via analogy is central to humans’ remarkable capacity to generalize knowledge to novel situations (Carey, 2011, Hofstadter, 2001, Lake et al., 2017). Many benchmarks have been designed to evaluate abstract reasoning abilities in machines (Foundalis, 2025, Hofstadter, 1995, Zhang et al., 2019). Among the most prominent such benchmarks is the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019). ARC consists of a set of idealized problems that require few-shot rule-induction and analogical reasoning. As Figure 1 shows, each puzzle (“task”) consists of a small set of demonstrations—initial and transformed grids—and a test grid, each ranging in size from $1\times 1$ to $30\times 30$, with each cell having one of 10 possible colors. To solve a task, an agent should infer a rule governing the demonstrations and apply that rule to the test input to produce a correct output grid.
@@
Chollet 2025 devised 1,000 such tasks, releasing 400 easier puzzles as a “training set,” 400 harder puzzles as an “evaluation set,” and keeping the remaining harder puzzles to form private test sets. Participants in the 2024 ARC-AGI Prize competition entered programs to vie for monetary prizes, including a \$600,000 grand prize for a program that exceeds 85% accuracy—that is, percentage of correct output grids—on a private test set of 100 tasks [^1] Average human performance was measured on the comparable, but somewhat easier public evaluation set as 64% (LeGris et al., 2024).. The top scoring program, which employed a fine-tuned LLM and extensive data augmentation, reached about 54% accuracy (Chollet et al., 2024).
@@
After the competition, Chollet and colleagues, with collaboration from OpenAI, tested a pre-release version of OpenAI’s o3 model on a different “semi-private” test set of 100 tasks. This model achieved 76% accuracy on its low-effort setting and 88% accuracy on its high-effort setting, with computing cost per task estimated at \$200 and \$20,000 respectively (Chollet et al., 2025). While o3-preview was not qualified to participate in the official competition, [^2] OpenAI’s o3 model violated two key competition rules: it could not be run locally on the competition server and its processing required substantially more than the required time limit. its superior performance was described as “a genuine breakthrough, marking a qualitative shift in AI capabilities compared to the prior limitations of LLMs”(Chollet, 2024).
@@
Despite the high accuracy of o3 on ARC tasks, it is not clear to what extent AI systems have achieved human-like abstract reasoning abilities. Consider the task illustrated in the top row of Figure 1. A human solving this task is likely to be able to generalize across different instantiations of the underlying abstract concepts—identifying and removing the top and bottom objects—no matter the size, shape, color, position, or number of objects. To our knowledge, no prior studies have assessed whether AI systems such as o3 are solving these tasks by using the intended, generalizable abstractions, or if they are inferring less generalizable rules (“shortcuts”) based on unintended correlations in task demonstrations.
@@
Here we assess the abstractions used by several commercial and open-weight models in solving tasks from ConceptARC (Moskvichev et al., 2023), a benchmark in the ARC domain containing tasks organized around basic spatial and semantic concepts, such as “inside vs. outside,” “above vs. below,” “extend to boundary,” and “same vs. different.” For example, the tasks shown in Figure 1 are from ConceptARC’s “top vs. bottom” and “extract object” concept groups, respectively. As described in Moskvichev et al. (2023), ConceptARC was designed to test robust understanding of these concepts by providing tasks—designed to be simple for humans—that deploy each concept in varying contexts and require varying degrees of generalization. Because it isolates simple abstract concepts, we believe this benchmark to be better suited than the original ARC dataset for investigating the concepts used by humans or machines in solving tasks.
@@
Figure 1: Each row shows a task from the ConceptARC benchmark. Each task shown consists of three demonstrations of a transformation and one test grid. In this study, the solver is tasked with generating a rule that describes the transformations and applying that rule to the test grid.
@@
Previous evaluations using the o3 model (as well as all entries in the 2024 ARC-AGI Prize competition) relied on text-based representations of the demonstration and test grids to solve each ARC task. Each grid is represented as an integer matrix, with entries encoding colors indexed from 0 to 9. However, o3 and related models are reported to possess sophisticated reasoning abilities in both textual and visual modalities (OpenAI, 2025). In our experiments, we investigate the models’ abstract reasoning abilities in both modalities. We also examine how reasoning effort (the token budget allocated for the reasoning stage) and access to external “tools” (here, the ability to generate and execute Python code) affect a model’s ability to discover abstract rules and solve tasks.
@@
In the following sections, we describe our experimental setup and results, and discuss how our findings relate to three central questions: (1) How does the accuracy achieved by AI models on ConceptARC tasks compare to that of humans? (2) To what extent do the rules generated by AI models and by humans capture the abstractions intended by the test designers, and to what extent do they rely on unintended, superficial patterns? (3) How do modality (textual vs. visual), reasoning effort (token budget), and Python tool access affect how well models can solve these tasks via the intended abstractions?
@@ 2.1 Dataset
To create ConceptARC, Moskvichev et al. 2023 chose 16 basic spatial and semantic concepts, and for each concept created 30 tasks that focused on that concept in different instantiations, with different degrees of abstraction, for a total of 480 tasks. Unlike the original ARC corpus, all ConceptARC tasks were designed to be relatively easy for humans, since each relies on a simple abstract concept and a straightforward application of that concept to a novel test grid.
2.2 Experiments on AI Models
@@
We evaluated four proprietary multimodal “reasoning” models on the ConceptARC tasks: OpenAI’s o3 and o4-mini, Google’s Gemini 2.5 Pro, and Anthropic’s Claude Sonnet 4. For comparison, we also evaluated three non-reasoning multimodal models: OpenAI’s GPT-4o, Meta’s Llama 4 Scout, and Alibaba’s Qwen 2.5 VL 72B. To maximize reproducibility, non-reasoning models were run with temperature 0. Because the o3, o4-mini, and Claude Sonnet 4 APIs restrict temperature to 1.0, we used temperature 1.0 for all four reasoning models to maintain comparability. For experiments using the textual modality, we used the same prompt as in Chollet et al.’s 2024 evaluation of o3-preview. For experiments using the visual modality, we used a slightly modified version of this prompt. These prompts are given in full in Appendix A and Appendix B.
@@
For both modalities, models were asked to generate a JSON object containing the transformation rule and the corresponding output grid, represented as a matrix of integers. This setup enables two-fold evaluation: (i) grid output accuracy and (ii) the degree to which model-generated rules capture the tasks’ intended abstractions. We evaluated human-generated solutions to these tasks using the same criteria, analyzing unpublished data [^3] A. Moskvichev, personal correspondence obtained from the study reported by Moskvichev et al. 2023 in which humans (participants on the Prolific Academic platform) were presented with ConceptARC tasks as images and asked to produce both the correct output grids and the rules they used to generate them. For each model setting, each task was given in an independent prompt, with the context window reset (cleared) before a new task was given. Due to resource constraints, we report pass@1 results for both AI models and humans. [^4] The ARC Prize competition reported pass@2 results, that is, two independent runs on each task. If one of the runs produces a correct output grid, the task is counted as correctly solved. Moskvichev et al. 2023 reported pass@3 results for both humans and the GPT models they tested.
@@
We evaluated o3 under its low- and medium-effort reasoning settings. [^5] OpenAI does not specify the token budget allocated to these settings. Due to resource constraints, we did not test the high-effort setting. We evaluated Gemini 2.5 Pro and Claude Sonnet 4 with a reasoning budget of 16,000 tokens, chosen to roughly approximate OpenAI’s medium-effort setting. Additionally, across all reasoning models and modalities, we evaluated two tool-access conditions: one in which Python tools were enabled and one in which they were not.
2.3 Evaluating Responses of Humans and AI Models
@@
Evaluating output-grid accuracy in human and model responses is straightforward, since each task’s ground-truth solution is given in the ConceptARC corpus. For each task, humans were given the demonstrations and test grid images and were asked to generate the output grid using a custom editing tool, and models were asked to generate the output grid as a matrix with colors encoded as integers. The resulting grids can be compared automatically with the ground truth; only exact matches are considered correct.
@@
While output-grid accuracy has been widely used to assess AI model performance on ARC tasks, to our knowledge, no prior studies have investigated whether output-grid correctness on ARC tasks reflects a grasp of the intended abstract concepts underlying the tasks, or alternatively, the extent to which correctness can be achieved by identifying and exploiting unintended, superficial patterns (“shortcuts"). It is well known that large neural-network models are capable of discovering spurious patterns in data and using these patterns to arrive at correct answers (Du et al., 2023, Geirhos et al., 2020).
To address this question, in our experiments we asked the models to output not only the transformed test grid but also a natural-language rule describing the transformation. Moskvichev et al. 2023 similarly collected such rules from their human participants, though only for correctly solved tasks.
Evaluating correctness of natural-language rules requires human judgment. [^6] We experimented with using LLMs to automatically evaluate rule correctness, but this was not sufficiently successful; further investigating LLM capabilities for this task would be interesting for future work. Accordingly, we manually annotated both model- and human-generated rules on their quality, distinguishing between “incorrect” (rules that do not work on the demonstrations, as well as rare cases where the model did not return a rule), “correct-unintended” (rules that work on the demonstrations, but do not capture the intended abstractions), and “correct-intended” (rules that align with the intended abstractions).
For example, in the first task shown in Figure 1, a sample human-generated rule is “You will see a number of shapes in a vertical orientation. Copy the grid from the example then remove the topmost and bottom-most shapes. The middle shapes remain unchanged.” We rated this rule as correct-intended. o3’s generated rule, from our experiment with textual inputs and medium reasoning effort, was “remove any object that intersects either the topmost or bottommost non-empty row, replacing it with zeros,” which we also rated as correct-intended. o3’s generated rule in our experiment with visual input and medium reasoning effort was “Delete the highest and lowest coloured components in the grid (the first and last contiguous non-black groups when scanning from top to bottom); keep the rest unchanged.” Also correct-intended, even though the output grid was incorrect in this case.
In the second task in Figure 1, a sample human-generated rule is “Take the shape that is imposed onto the background pattern and simply use that shape as the solution.” We rated this as correct-intended. o3 with textual input and medium effort generated the rule “Find the single colour absent from both the first row and the first column. Crop the minimal rectangle that contains all occurrences of this colour, keep that colour and set everything else inside the crop to 0.” We rated this as correct-unintended, since it correctly described the three demonstrations, but ignored the intended abstraction of extracting an object. In fact, we noticed that, given the textual representation of a task, o3 most often phrased its rules in terms of colors or individual pixels rather than objects. o3 with visual input and medium effort generated the rule “Crop the minimal bounding box around the unique 1-cell-thick closed loop. Leave the loop’s colour unchanged and recolour every other cell in the crop with the most frequent colour of the original grid,” which we rated as “incorrect”.
While it is not certain that the natural-language rule a model generates for a given task is a faithful description of how it solved that task, we did find that, at least in the textual setting, the generated rules did tend to align with the generated output grid, supporting the hypothesis that these rules were generally faithful.

### v4

@@ Abstract
OpenAI’s o3-preview reasoning model exceeded human accuracy on the ARC-AGI-1 benchmark, but does that mean state-of-the-art models recognize and reason with the abstractions the benchmark was designed to test? Here we investigate abstraction abilities of AI models using the closely related but simpler ConceptARC benchmark. Our evaluations vary input modality (textual vs. visual), use of external Python tools, and reasoning effort. Beyond output accuracy, we evaluate the natural-language rules that models generate to explain their solutions, enabling us to assess whether models recognize the abstractions that ConceptARC was designed to elicit. We show that the best models’ rules are frequently based on surface-level “shortcuts,” capturing intended abstractions considerably less often than humans. In the visual modality, AI models’ output accuracy drops sharply; however, our rule-level analysis reveals that a substantial share of their rules capture the intended abstractions, even as the models struggle to apply these concepts to generate correct solutions. In short, we show that using accuracy alone to evaluate abstract reasoning can substantially overestimate AI capabilities in textual modalities and underestimate it in visual modalities. Our results offer a more faithful picture of AI models’ abstract reasoning abilities and a more principled way to track progress toward human-like, abstraction-centered intelligence.
ARC, Multimodal Reasoning, Abstraction, Vision Language Models
@@ 1 Introduction
The ability to quickly form abstractions and use them to reason by analogy is central to humans’ remarkable capacity to generalize knowledge to novel situations (Carey, 2011; Hofstadter, 2001; Lake et al., 2017). Many benchmarks have been designed to evaluate abstract reasoning abilities in machines (Foundalis, 2025; Hofstadter, 1995; Zhang et al., 2019). Among the most prominent is the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019), [^1] In 2023 ARC was renamed “ARC-AGI,” but for simplicity we will refer to it here as “ARC.” ARC consists of a set of idealized problems that require few-shot rule-induction and analogical reasoning. As Figure 1 shows, each puzzle (“task”) consists of a small set of demonstrations (initial and transformed grids) and a test grid, each ranging in size from $1\times 1$ to $30\times 30$, with each cell having one of 10 possible colors. To solve a task, an agent must infer a rule governing the demonstrations and apply that rule to the test input to produce a correct output grid. According to ARC’s creator, François Chollet (2019), ARC is “built upon an explicit set of priors designed to be as close as possible to innate human priors. We argue that ARC can be used to measure a human-like form of general fluid intelligence and that it enables fair general intelligence comparisons between AI systems and humans.”
@@
Chollet (2025) created 1,000 such tasks, releasing 400 easier puzzles as a “training set” and 400 harder puzzles as an “evaluation set.” The remaining, more difficult puzzles were held out as private test sets. Participants in the 2024 ARC-AGI Prize competition entered programs that vied to achieve high accuracy (i.e., a high percentage of correct output grids) on a private test set of 100 tasks, with a \$600,000 grand prize for any program that exceeded 85% accuracy. [^2] Average human performance was measured at 64% on the comparable but somewhat easier public evaluation set (LeGris et al., 2025). The highest scoring program, with about 54% accuracy, employed a fine-tuned LLM, extensive data augmentation, and test-time training (Chollet et al., 2024).
@@
Following the competition, Chollet and colleagues tested a pre-release version of OpenAI’s o3 model on a different “semi-private” test set of 100 tasks. This model achieved an accuracy of 76% on its low-effort setting and 88% on its high-effort setting, with computing cost per task estimated at \$200 and \$20,000 respectively (Chollet et al., 2025). While o3-preview was not qualified to participate in the official competition, its performance was described as “a genuine breakthrough, marking a qualitative shift in AI capabilities compared to the prior limitations of LLMs” (Chollet, 2024).
@@
However, despite the high accuracy of o3 on ARC tasks, it is not clear whether AI systems have achieved the human-like abstract reasoning abilities that ARC was designed to test. Consider, for example, the task illustrated in the top row of Figure 1 (from the ConceptARC benchmark, described below). A human solving this task could likely generalize across different instantiations of the underlying abstract concepts, that is, identifying and removing the top and bottom objects, regardless of the size, shape, color, position, or number of objects. Yet to our knowledge, no prior studies have assessed whether AI systems such as o3 are solving these tasks by using the intended, generalizable abstractions, or whether they exploit less generalizable rules (“shortcuts”) based on unintended correlations in the demonstrations.
@@
Here we assess the abstractions used by several commercial and open-weight models in solving tasks from ConceptARC (Moskvichev et al., 2023), a benchmark in the ARC domain containing tasks organized around the same “core knowledge priors” that the ARC benchmark was designed to test. ConceptARC’s tasks are focused on concepts such as “inside and outside,” “above and below,” or “same vs. different.” For example, the tasks shown in Figure 1 are from ConceptARC’s “top vs. bottom” and “center” concept groups, respectively. As described in Moskvichev et al. (2023), ConceptARC was designed to test robust understanding of these concepts by providing tasks (designed to be simple for humans) that use each concept in various contexts and require varying degrees of generalization to solve. As ConceptARC isolates simple abstract concepts, we believe this benchmark is better suited than ARC for investigating the concepts used by humans and machines in solving tasks. ARC is substantially harder for humans than ConceptARC (Moskvichev et al., 2023), and frequently employs compositional reasoning, which would make it necessary to disentangle different types of intended abstractions, complicating the evaluation process.
@@
Figure 1: Each row shows a task from the ConceptARC benchmark. A task consists of three demonstrations of a transformation and one test grid. In this study, the solver is tasked with generating a rule that describes the transformations and applying that rule to the test grid.
@@
Previous evaluations using the o3 model, as well as all entries in the 2024 ARC-AGI Prize competition, relied on the same text-based representation of the demonstration and test grids to solve each ARC task. Each grid is represented as an integer matrix, with entries encoding colors indexed from 0 to 9. However, because o3 and related models are reported to possess sophisticated reasoning abilities in both textual and visual modalities (OpenAI, 2025), we investigate the models’ abstract reasoning abilities in both modalities. We also examine how reasoning effort (the token budget allocated for the reasoning stage) and access to external “tools” (here, the ability to generate and execute Python code) affect a model’s ability to discover abstract rules and solve tasks.
@@
In the following sections, we describe our experimental methodology and results, and discuss how our findings relate to three central questions: (1) How does the accuracy achieved by AI models on ConceptARC tasks compare to that of humans? (2) To what extent do the rules generated by AI models and by humans capture the abstractions intended by the test designers, and to what extent do they rely on unintended, superficial patterns? (3) How do modality (textual vs. visual), reasoning effort (token budget), and Python tool access affect how well models can solve these tasks via the intended abstractions?
@@ 2 Methodology
To create ConceptARC, Moskvichev et al. (2023) chose 16 basic spatial and semantic concepts, and for each concept created 30 tasks that focused on that concept in different instantiations, with different degrees of abstraction, for a total of 480 tasks. In contrast to ARC tasks, all ConceptARC tasks were designed to be relatively easy for humans to solve: each focuses on a simple abstract concept and its application to a test grid. Moskvichev et al. (2023) reported human accuracy on ConceptARC to be 91% (based on three guesses per task), which can be roughly compared to 64% human accuracy on the ARC evaluation set reported in LeGris et al. (2025) (based on two guesses per task).
@@
We evaluated four proprietary multimodal “reasoning” models on the ConceptARC tasks: OpenAI’s o3 and o4-mini, Google’s Gemini 2.5 Pro, and Anthropic’s Claude Sonnet 4. For comparison, we also evaluated three non-reasoning multimodal models: OpenAI’s GPT-4o, Meta’s Llama 4 Scout, and Alibaba’s Qwen 2.5 VL 72B. To maximize reproducibility, non-reasoning models were run with temperature 0. Because the APIs for o3, o4-mini, and Claude Sonnet 4 restrict temperature to 1.0, we used temperature 1.0 for all four reasoning models to maintain comparability. For experiments using the textual modality, we used the same prompt as in Chollet et al.’s (2024) evaluation of o3-preview. For experiments using the visual modality, we used a slightly modified version of this prompt. Full prompts are given in Appendices A, B, and Appendix C.
@@
For both modalities, models were asked to generate a JSON object containing the transformation rule and the corresponding output grid, represented as a matrix of integers (the same representation used in previous ARC evaluations). This setup enables two-fold evaluation: (i) grid output accuracy and (ii) the degree to which model-generated rules capture the tasks’ intended abstractions. We evaluated human-generated solutions using the same criteria, analyzing unpublished data obtained from the study reported by Moskvichev et al. (2023) in which humans (participants on the Prolific Academic platform) were presented with ConceptARC tasks as images and asked to produce both the correct output grids and the rules they used to generate them. [^3] Unpublished data was provided by A. Moskvichev, personal correspondence. For each model setting, each task was given in an independent prompt, without prior context. Due to resource constraints, we report pass@1 results for both AI models and humans. [^4] The ARC Prize competition reported pass@2 results; that is, two independent runs on each task. If one of the runs produces a correct output grid, the task is counted as correctly solved.
@@
We evaluated o3 under its low- and medium-effort reasoning settings. [^5] OpenAI does not specify the token budget allocated to these settings. Due to resource constraints, we did not test the high-effort setting. We evaluated Gemini 2.5 Pro and Claude Sonnet 4 with a reasoning budget of 16,000 tokens, which roughly approximates OpenAI’s medium-effort setting. Additionally, for reasoning models we evaluated two tool-access conditions: one in which Python tools were enabled and one in which they were not.
Evaluating Responses of Humans and AI Models
@@
For each task, humans were given the demonstrations and test grid images and were asked to generate the output grid using a custom editing tool, whereas models were asked to generate the output grid as a matrix with colors encoded as integers. Following the ARC evaluation criteria, to be considered correct, a generated output-grid must exactly match the ground-truth solution in the ConceptARC corpus, and be in the requested format. For a more detailed analysis of format deviations see Appendix K.
@@
While output-grid accuracy has been widely used to assess performance on ARC tasks, to our knowledge, no prior studies have investigated the extent to which accuracy actually reflects a grasp of the intended abstract concepts underlying the tasks versus the identification of unintended, superficial patterns (“shortcuts”) in the data. It is well known that large neural-network models are capable of discovering spurious patterns in data and using these patterns to arrive at correct answers (Du et al., 2023; Geirhos et al., 2020). To investigate the extent to which AI models are using human-like abstractions to solve tasks, we asked the models to output not only the transformed test grid but also a natural-language rule describing the transformation. Moskvichev et al. (2023) similarly collected such rules from their human participants, though only for correctly solved tasks.
Evaluating the correctness of natural-language rules requires human judgment. Accordingly, we manually classified both model- and human-generated rules using three possible classes: “correct-intended” (rules that align with the intended abstractions), “correct-unintended” (rules that work on the demonstrations but do not capture the intended abstractions), and “incorrect” (rules that do not work on the demonstrations). [^6] In cases where the model or human participant did not provide a rule, or the rule was too unclear to confidently assess, the rule was annotated as “non-responsive” or “unclear,” respectively. For example, in the first task shown in Figure 1, a sample human-generated rule is “Use the same grid size as in the input. There will be multiple objects. Simply remove the top object and the bottom object only.” We rated this rule as correct-intended. Claude Sonnet 4’s generated rule, from our experiment with textual inputs and medium reasoning effort, was “Remove the topmost and bottommost colored regions, keep all colored regions in between,” which we also rated as correct-intended. Gemini 2.5 Pro’s rule in this same setting was “Identify all connected shapes of non-zero numbers. Sort these shapes first by their size (number of cells) in ascending order, then by their color value in ascending order as a tie-breaker. Remove the two shapes that come first in this sorted list, and keep all other shapes.” Here, Gemini is using the ordering of integer codes for colors, which is meant to be an arbitrary textual encoding, as a relevant feature. It turns out that Gemini found an unintended but correct pattern in the demonstrations, so this rule was classified as correct-unintended, even though Gemini’s generated output grid, which followed its rule, was incorrect. Finally, o3’s generated rule in this same setting was “Remove all smallest objects: find each connected non-zero color component, count its cells, identify the minimum area, and delete every component having that minimum size (replace its cells with 0); leave every other cell unchanged.” As this rule did not correctly describe the demonstrations, we classified it as incorrect.
In the second task in Figure 1, a sample human-generated rule is “Select the colour that is in the middle of the input. Only output that colour as a $1\times 1$ grid.” As before, we rated this as correct-intended. o3 with textual input, medium effort, and tool use enabled, generated the rule “Return the middle element of the 1-D input list (index $\lfloor n/2\rfloor$ for odd length) as a single-value output grid.” Also correct-intended. Claude’s rule in the same setting was “Find all values that appear with the minimum frequency in the input grid, sort them in ascending order, then return the second smallest value if multiple values exist, or the only value if just one exists.” This rule again uses the specific integer values encoding colors, and is a correct, though unintended, description of the demonstrations, and thus was classified as correct-unintended; it also happens to generate a correct output grid. In the same setting, Gemini generated a very similar correct-unintended rule that also generated a correct output grid.
Further examples of correct-unintended rules are given in Appendix D. It is important to clarify that we do not consider “correct-unintended” rules to be incorrect; rather, they reflect a mismatch between the generated rule and the one intended by the task designer, which is based on humanlike “core knowledge” concepts (Chollet, 2019) that both ARC and ConceptARC were designed to evaluate.
While it is not guaranteed that a model’s natural-language rule for a given task is a faithful reflection of how the task was solved, we manually analyzed the alignment between generated rules and their corresponding output grids. We found that, across all experimental settings, the output grid, whether correct or not, was faithful to the model’s rule in over 90% of the cases we examined, providing evidence that these rules serve as reasonable proxies for the model’s reasoning process. Further details are provided in the next section.